OpenStudy (anonymous):

hi there! ok unsure about my answer on this one. . . 9. Calculate the experimental value for density of the aluminum shot based on its mass (19.95) as well as its volume (8), using the following formula: Density of Al = mass of Al shot/ volume of Al shot

OpenStudy (anonymous):

So Density of Al = 19.95 - 8

jimthompson5910 (jim_thompson5910):

I think you meant to say Density of Al = 19.95/8

OpenStudy (anonymous):

oops yes i did

jimthompson5910 (jim_thompson5910):

divide that out to get your answer

OpenStudy (anonymous):

im not understanding what to do after finding D = 2.5. . . .

OpenStudy (anonymous):

is that the experimental value in itself?

jimthompson5910 (jim_thompson5910):

that's the experimental density assuming 19.95 is the experimental mass and/or 8 is the experimental volume

jimthompson5910 (jim_thompson5910):

if both are exact values, then it's not experimental (since nothing in an experiment is perfect)

OpenStudy (anonymous):

thats why i was a bit confused. .. coz then i plug it into the next step of the equation which is: 10. If the accepted value for the density of aluminum is 2.70 g/mL, calculate the percent error of your density calculation, using the formula listed in #5 above. /2.70 – 2.5/ * 100/ 2.70 = .2/2.70 = .074 X 100 = 7.4& error of calculation

jimthompson5910 (jim_thompson5910):

you should have gotten 2.49375 after you divide 19.95/8

OpenStudy (anonymous):

wouldnt i round to 2.5?

jimthompson5910 (jim_thompson5910):

|2.49375 - 2.7| --------------- * 100 2.7 |-0.20625| --------------- * 100 2.7 0.20625 --------------- * 100 2.7 0.07638888888889* 100 7.638888888889 %

jimthompson5910 (jim_thompson5910):

So the error is roughly 7.638888888889 %

jimthompson5910 (jim_thompson5910):

you could round to make things shorter to get 7.64%

OpenStudy (anonymous):

where am i going wrong in my calculations? how did i end up with 7.41

jimthompson5910 (jim_thompson5910):

you used 2.5 instead of 2.49375

jimthompson5910 (jim_thompson5910):

so that explains the slight difference

OpenStudy (anonymous):

thank you so much for your patience!

jimthompson5910 (jim_thompson5910):

you're welcome